OpenStudy (anonymous):

PLEASE HELP: Suppose that in 2001, the average cost of a gallon of milk was $2.43. In 2007 the average price was $4.21. Assuming the relationship is linear, find the equation of the line that models the price of milk over time since 1999. Round to the nearest thousandth.

OpenStudy (anonymous):

x is the year and y is the price general linear exquation: y=kx+d i. 2.43=k*2001+d ii. 4.21=k*2007+d now make the gaussian elimination i-ii and you get your equation. then put 1999 in x and you have the result

OpenStudy (anonymous):

can you explain more im confused

OpenStudy (anonymous):

I'll answer it in an hour if nobody else does till then. Can't atm